OpenStudy (anonymous):

Step By Step simplify √(72)/(25).

OpenStudy (anonymous):

Allthese problems u posted, do same way, factor the numbers under the square root eg 72= 2*2*2*3*3 and u can take the 3*3 "outside" the sqrt as 3 25 = 5*5 so u can take out as 5.